Two point charges 20μC and 10μC are separated by 0.05m in free space.Find the force between them.Also calculate the force between them. Also calculate the force when a dielectric medium of dielectric constant 3 is introduced between them.

Grade:12

1 Answers

Khimraj
5 years ago
initial force
F = Kq1*q2/r2 = 9*109*20*10-6*10*10-6/25*10-4 = 720 N
final force will be K times the initial force = 3*720 = 2160N
Hope it clears.....................................